3x+1=7x+17
x= -4
But I don't understand the steps.
Thank you!

Found 2 solutions by NancyLam, Alan3354:
Answer by NancyLam(35)

3x+1=7x+17
x= -4
7x-3x=1-17
4x=-16
x=-16/4
x=-4

Answer by Alan3354(69443)

3x+1=7x+17
Subtract 1 from both sides
3x = 7x + 16
Subtract 3x from both sides
0 = 4x + 16
Subtract 16 from both sides
-16 = 4x
Divide by 4
x = -4
---------
The steps can be in a different order, and in different ways, but the result is the same.
3x+1=7x+17
Subtract 17 from both sides
3x - 16 = 7x
Subtract 3x
-16 = 4x
x = -4